# Magnetic field measurements in a sample of Class I and Flat-Spectrum protostars observed with SPIRou

### 18 Mar 2026 v1

@@ Appendix E LSD Profiles and parameter values
Figure 8: LSD profiles for each star. Different colours are used for the different observations. The Y-axes are II/IcIc, NN/IcIc, and VV/IcIc, where $Ic$ is the intensity of the stellar continuum. The II profile represents the total intensity, the Stokes $V$ profile corresponds to the circular polarisation, and the Stokes $N$ profile corresponds to the null spectra. Pixel size of 4.8 instead of 2.4 is applied for stars with $v\sin i$ ¿ 35 km s-1.
Table 5: LSD Stokes II fitting results, veiling values, and longitudinal magnetic field detection and measurements [^11] The depth, the projected rotational velocity ($v\sin i$), the radial velocity ($v_{\rm rad}$), computed with a physical stellar model (see Appendix A). The veiling values in H and K band, the longitudinal magnetic field values, their corresponding uncertainties, the false alarm probability, and the magnetic detection status (defined as $D$ for definite detection, MD for marginal detection, and ND for no detection) are also presented. The stars indicated in bold are detected magnetic stars (at least one $D$ in the different observations). Sources with an * symbol are Flat-Spectrum targets.

### 22 Mar 2026 v2

@@ Appendix E LSD Profiles and parameter values
Figure 8: LSD profiles for each star. The different colours indicate the different observations. The Y-axes are II/IcIc, NN/IcIc, and VV/IcIc, where $Ic$ is the intensity of the stellar continuum. The II profile represents the total intensity, the Stokes $V$ profile corresponds to the circular polarisation, and the Stokes $N$ profile corresponds to the null spectra. A pixel size of 4.8 instead of 2.4 is applied for stars with $v\sin i$ ¿ 35 km s-1.
Table 5: LSD Stokes II fitting results, veiling values, and longitudinal magnetic field detection and measurements. [^11] From left to right the columns show the depth, the projected rotational velocity ($v\sin i$), the radial velocity ($v_{\rm rad}$), computed with a physical stellar model (see Appendix A), the false alarm probability, the magnetic detection status (defined as $D$ for definite detection, MD for marginal detection, and ND for no detection), the veiling values in H and K band, the longitudinal magnetic field values, and their corresponding uncertainties. The stars indicated in bold are detected magnetic stars (at least one $D$ in the different observations). Sources with an * symbol are flat-spectrum targets.